# MLP-SLAM: Multilayer Perceptron-Based Simultaneous Localization and Mapping With a Dynamic and Static Object Discriminator

### v1

@@ Abstract
The Visual Simultaneous Localization and Mapping (V-SLAM) system has seen significant development in recent years, demonstrating high precision in environments with limited dynamic objects. However, their performance significantly deteriorates when deployed in settings with a higher presence of movable objects, such as environments with pedestrians, cars, and buses, which are common in outdoor scenes. To address this issue, we propose a Multilayer Perceptron (MLP)-based real-time stereo SLAM system that leverages complete geometry information to avoid information loss. Moreover, there is currently no publicly available dataset for directly evaluating the effectiveness of dynamic and static feature classification methods, and to bridge this gap, we have created a publicly available dataset containing over 50,000 feature points. Experimental results demonstrate that our MLP-based dynamic and static feature point discriminator has achieved superior performance compared to other methods on this dataset. Furthermore, the MLP-based real-time stereo SLAM system has shown the highest average precision and fastest speed on the outdoor KITTI tracking datasets compared to other dynamic SLAM systems.The open-source code and datasets are available at https://github.com/TaozheLi/MLP-SLAM.
@@ I INTRODUCTION
Visual Simultaneous Localization and Mapping (V-SLAM) is designed to achieve precise ego-localization and construct an accurate map of the surroundings using only the camera sensor. The camera sensor is light-weight and cost-effective, while capable of capturing high resolution images. V-SLAM has played a crucial role in various fields, including robotics, space exploration, and autonomous driving. For example, accurate mapping and precise ego-localization significantly impact the navigation and decision-making processes of autonomous driving systems. However, most existing SLAM systems are heavily reliant on the assumption of a static world, including state-of-the-art (SOTA) systems such as ORB-SLAM2 [1], ORB-SLAM3 [2], LSD-SLAM [3] and Direct Sparse Odometry [4]. These systems typically perform with high precision in environments with predominantly static objects. Nevertheless, their precision deteriorates when deployed in dynamic environments, which are common in real world scenarios. Research efforts aimed at addressing the degradation of SLAM systems in dynamic environments can be categorized into three parts. Part I focuses only on geometry methods, which involve classifying feature points based solely on geometric information. This information includes epipolar error, reprojection error, image intensity, depth, and optical flow, as demonstrated in works such as [5]. Part II, exemplified by works like DynaSLAM [6], leverages prior class information obtained from object detection models or semantic segmentation models to identify features within bounding boxes or masks of objects belonging to dynamic classes as outliers, which are then discarded during camera pose estimation. While these approaches partially mitigate the degradation of SLAM systems in dynamic environments, their effectiveness is limited. For instance, features extracted from stationary parked cars on the roadside may be erroneously removed, despite their static nature and potential value in estimating camera pose, leading to information loss. In Figure 1, a qualitative comparison is presented between our Multilayer Perceptron (MLP)-based discriminator for dynamic and static feature points and the method proposed in DynaSLAM, using a snapshot from the Kitti tracking dataset [7]. The top image displays the results of our method, while the bottom image shows the results of the method in DynaSLAM. In the left image, the method in DynaSLAM categorizes feature points on the highway guardrail as dynamic, whereas our method correctly predicts them as static. In the right image, numerous static feature points associated with parked cars on the roadside are misclassified as dynamic by the method in DynaSLAM, whereas our method correctly identifies them as static. This evidence supports the effectiveness and reliability of our method. Part III represents the current mainstream approach, which combines methods from Part I and Part II, as demonstrated in works such as DynaSLAM, CFP-SLAM [8], PointSLOT [9], VDO-SLAM [10], and D-SLAM [11], by initially acquiring prior class information through deep learning models such as object detection models or semantic segmentation models. The most prevalent deep learning models utilized in the Simultaneous Localization and Mapping (SLAM) domain are from the YOLO series, known for their high precision and rapid processing speed, such as YOLOv5 [12] and YOLOACT [13]. Subsequently, leveraging the prior class information obtained from these models, geometric methods are employed to eliminate dynamic feature points. These geometric aspects encompass various factors, including epipolar error, re-projection error, image intensity, depth, and velocity (if tracking dynamic object pose). However, a significant portion of the existing methodologies in Part III only utilize partial geometric information. For instance, some methods solely focus on epipolar error [14], re-projection error [9], or a combination thereof [8]. These approaches are commonly employed in the dynamic SLAM field. Nevertheless, these methods often require manual parameter tuning, lack generality, and are excessively laborious to fine-tune. Furthermore, the evaluation of methods distinguishing between dynamic and static feature points is typically conducted indirectly through SLAM field metrics, as there is a lack of publicly available datasets specifically designed for this classification task. This indirect evaluation method tends to be inaccurate and unreliable due to the randomness in the Random Sample Consensus (RANSAC) process in camera pose estimation within the SLAM system.
@@
To address the aforementioned gaps, this letter introduces a MLP-based real-time stereo SLAM system and establishes publicly available datasets to classify dynamic and static feature points. The contributions outlined in this work are as follows:
@@
  • We introduce an open-source MLP-based real-time stereo system designed to enhance performance in dynamic environments. This system leverages comprehensive geometry information and eliminates the need for manual parameter configuration. To the best of our knowledge, it is the first time that a MLP machine learning method is applied to SLAM system as a dynamic and static feature points discriminator.
@@
  • We have constructed publicly available datasets comprising more than 50,000 feature points sourced from the KITTI tracking datasets [7] for the purpose of training and testing dynamic and static feature points discriminators. The efficacy of it can be evaluated directly with this dataset.
@@ II-A Dynamic SLAM
The emergence of DynaSLAM has garnered significant attention in the Visual Simultaneous Localization and Mapping field since its release. This approach identifies feature points located within the semantic mask of potential moving objects as outliers and subsequently discards them. While it has shown improvements in the KITTI tracking sequences 01, 08, and 09 compared to ORB-SLAM2, this method may misrepresent dynamic and static objects, as shown in Figure 1. Moreover, DynaSLAM is plagued by high latency issues attributed to its reliance on the computationally intensive deep learning model MaskRCNN [15] for generating semantic segmentation masks. Subsequent research endeavors have aimed to address the degradation of SLAM systems in dynamic environments. For instance, work CFP-SLAM[9] introduced a Moving Objects Recognition (MOR) approach to differentiate between dynamic and static feature points by utilizing the re-projection error. In comparison, work [8] employs a DBSCAN[16] clustering machine learning technique to filter out features within bounding boxes that belong to the background and then employ a geometric method that incorporates geometric information such as re-projection error and epipolar error to distinguish dynamic and static feature points. An inherent limitation in [9] and [8] is the requirement of the establishment and tuning of manual parameters, which restricts the application scenarios and diminishes the generality of the SLAM system. Furthermore, these approaches do not fully leverage geometric information. The VDO-SLAM[10] and DynaSLAM II[17] are are extensions of DynaSLAM, with DynaSLAM II representing the current state-of-the-art (SOTA) in camera motion estimation. Both works initially estimate the camera pose using feature points outside the potential dynamic class semantic mask, akin to the approach in DynaSLAM. Subsequently, they estimate the object pose for each frame and optimize the camera and object poses simultaneously through Bundle Adjustment (BA) factor graph. However, these methodologies encounter high latency issues due to data association challenges between dynamic objects and object pose estimation. Additionally, the precision of object pose estimation is compromised by the limited number of matched feature points and the quality of feature matching for each object.
@@ II-B Multilayer Perceptron
Multilayer Perceptron (MLP) [18] is commonly utilized as a tool for classification tasks, and it has been effectively employed in artificial intelligence, particularly in the field of computer vision. One of the most renowned applications is the utilization of MLP to address the problem of recognizing handwritten digits [19]. Another application is the diagnosis systems in the medical filed [20]. The MLP primarily comprises three components. The first component is the input layer, which is responsible for receiving the features of each data point. The second component is the hidden layer, where the number of nodes in each layer and the depth of the hidden layer significantly influence its performance. Specifically, the model can handle more complex tasks with a greater depth in the hidden layer and more nodes in each hidden layer. The third component is the output layer, which can be viewed as the post-processing of results from the hidden layers to fulfill various tasks. In the field of Simultaneous Localization and Mapping (SLAM), studies [21, 5, 22] applied the MLP to decode grid features into occupancy values in the mapping part. To the best of our knowledge, the method introduced in our letter is the first instance of applying the MLP as a discriminator for dynamic and static feature points in the SLAM domain.
@@ III-A Notations in SLAM
World and camera coordinates are denoted by {$W$} and {$C$} respectively. ${}^{t}X^{B}_{i}$ describes the $i_{\text{th}}$ 3D cloud point at time $t$ and located in the coordinate of {$B$}, $B\in\{W,C\}$. Moreover, ${}^{t}P_{i}$ marks the $i_{\text{th}}$ point in the image plane, and the operation $\pi(.)$ represents projecting an 3D point cloud from camera coordinates into the image plane. i.e. ${}^{t}P_{i}=\pi(^{t}X^{C}_{i})$. Besides, the translation matrix ${}^{t}T_{CW}=\begin{bmatrix}{}^{t}R_{CW},\textbf{t}_{CW}\\ 0,1\end{bmatrix}\in SE(3)$ is used to describe a transformation from the world coordinate $\{W\}$ to the camera coordinate $\{C\}$. The matrix ${}^{t}R_{CW}\in SO(3)$ denotes the rotation part and the $\textbf{t}_{CW}$ denotes the translation part of ${}^{t}T_{CW}$. The essential matrix between two consecutive frames is denoted by $E$. The $i_{\text{th}}$ feature point is denoted by $x_{i}$, and its depth is represented by $d_{i}$.
@@ III-B Camera Pose Estimation
In the camera pose estimation process, we search the static feature point correspondences among consecutive frames to avoid the impact of dynamic objects. More specifically, we first do a coarse estimation with feature points that are outside of the bounding box which belongs to potential dynamic objects, as well as the static feature points inside the bounding box that are identified by the Depth Filter Module. And then we do a fine estimation where static feature points on potential dynamic objects are filtered by the MLP model and they will also be used for camera pose estimation. The camera pose is acquired through minimizing the re-projection error of static feature point correspondences, which is described by:
@@
${}^{t}e_{i}={{}^{t}}P_{i}-\pi(^{t}T_{CW}{{{}^{t-1}}X^{W}_{i}}),$ (2)
and the operation $\rho(\cdot)$ is a robust kernel function for reducing impact of feature points with big error, and $N$ denotes the number of static feature correspondences.
@@ III-C Notations in MLP
$e_{I}(x_{1},x_{2})$ $=\left\lVert I(P_{1})-I(P_{2})\right\rVert^{2}_{2},$ (3)
$e_{D}(E,x_{1},x_{2})$ $=\frac{P_{2}^{T}E}{{\left\lVert P_{2}^{T}E\right\rVert{}}}\cdot P_{1},$ (4)
$e_{Re}(R_{12},t_{12},x_{1},x_{2})$ $=\left\lVert\pi(R_{12}\pi(P_{2},d_{2})^{-1}+t_{12})-P_{1}\right\rVert^{2}_{2}.$ (5)
We use operation $M(\cdot)$ to denote applying MLP model for classifying dynamic and static feature points. The prediction result and the label are denoted by $y^{p}\in R^{2}$ and $y^{t}\in R^{2}$. And the label of each feature point is denoted by $c_{i}$. For our problem, the label $c_{i}$ equals 00 or $1$. We used 00 to indicate static class and $1$ for dynamic class. Thus, the MLP model predicted process can be described by
$c_{i}=argmax(y_{i}^{p}),$ (6)
where
$y^{p}_{i}=M(v_{i}).$ (7)
III-D Loss function
There are lots of different types of loss function. They all operate by quantifying the difference between a model’s predictions and actual target. With the loss function, the gradient can be calculated to update model parameters through back propagation. In this letter, Binary Cross-Entropy (BCE) Loss is chosen as loss function for MLP to classify dynamic and static feature points. The entropy indicates a calculation of the degree of randomness or disorder within a system. The cross-entropy measures the differences between two probability distributions that can be used to identify an observation. The BCE Loss is used for measuring the difference between the prediction of machine learning algorithm and the actual target prediction that is calculated from the negative value of the summation of the logarithm value of the probabilities of the predictions made by the machine learning algorithm against the total number of data samples. We use $y_{t}$ and $y_{p}$ to present true target and prediction of Multilayer perceptron model respectively. The mathematical equation for Binary Cross-Entropy (BCE) Loss is presented as following:
$BCE(y^{t}_{i},y^{p}_{i})=-y^{t}_{i}*log(y^{p}_{i})-(1-y^{t}_{i})*log(1-y^{p}_{i})$ (8)
@@ IV Proposed Method
Figure 2: Diagram of our proposed MLP based real-time stereo SLAM system. There are four main parts in total. The blue section involves pre-processing, feature extraction, object detection, and object tracking. The green section is for coarse estimation, generating a rough pose estimation. The pink section involves an MLP model for classifying dynamic and static feature points within potential dynamic objects. The yellow section is for fine estimation, refining the camera pose generated in the coarse estimation stage by minimizing the re-projection error of all static feature points. These points include those outside the bounding box of potential dynamic objects, filtered by the Depth Filter Module, and classified as static by the MLP model.
The diagram of our MLP-based stereo real-time SLAM system is shown in Figure 2. Initially, the input stereo images undergo feature extraction. Subsequently, object detection and tracking are performed on the left image. The process then moves to a coarse estimation stage where features outside the bounding box of potential dynamic objects, and the background static features within the bounding box that are filtered by the Depth Filter Module, are used for coarse camera pose estimation. Then each feature point $x_{i}$ that is inside the bounding box and not filtered by the Depth Filter Module generates an input feature $v_{i}$ using formulas (3), (4), and (5). The MLP model assigns a label $c_{i}$ to each feature point to distinguish static and dynamic feature points. Finally, all static feature points are used for fine camera pose estimation. In this section, details of each module and our collected dynamic and static feature points classification dataset will be elaborated.
IV-A Depth Filter Module
We opt for object detection and tracking models over semantic segmentation models in our SLAM system to reduce latency, as the latter typically requires more computational resources. However, object detection models produce less accurate bounding boxes compared to mask from semantic segmentation models. To address this issue, we introduce the Depth Filter Module to remove background features within the bounding box, classifying them as static. The process of the Depth Filter Module is as follows: Initially, we calculate the average depth $\mu_{d}$ of all feature points within the bounding box, provided there are a sufficient number of feature points. Subsequently, we calculate the standard deviation $\delta_{d}$ of these points. Any feature points falling outside the range $[\mu_{d}-\eta\delta_{d},\mu_{d}+\eta\delta_{d}]$ are eliminated from the potential dynamic feature point set and treated as static in coarse estimation stage. Here, $\eta$ represents a constant value that describes the degree of deviation from the average depth value.

### v2

@@ Abstract
The Visual Simultaneous Localization and Mapping (V-SLAM) system has advanced significantly in recent years, achieving high precision in environments with few dynamic objects. However, its performance declines in settings with numerous movable objects, such as those with pedestrians, cars, and buses, which are common in outdoor scenes. While some researchers are tackling this issue with artificial mathematical methods, these often suffering less generality and information loss. To address this problem, we propose a novel Multilayer Perceptron (MLP)-based method that utilizes complete geometric information to prevent information loss and capable to apply in various situations. To best of our knowledge, this is the first time that machine learning-based method introduced for distinguishing dynamic and static feature points in the SLAM field. Additionally, there is currently no publicly available dataset for evaluating dynamic and static feature classification methods. To fill this gap, we have created a publicly available dataset containing over 50,000 feature points. Experimental results show that our MLP-based dynamic and static feature point discriminator outperforms other methods on this dataset. Furthermore, our proposed MLP-based real-time stereo SLAM system achieves the highest average precision and fastest speed on the outdoor KITTI Odometry and KITTI360 datasets compared to other dynamic SLAM systems. The open-source code and datasets will be accessible at https://github.com/TaozheLi/MLP-SLAM upon paper acceptance.
@@ I INTRODUCTION
Visual Simultaneous Localization and Mapping (V-SLAM) allows for precise ego-localization and accurate mapping using a lightweight, cost-effective camera sensor. It is vital in robotics, space exploration, and autonomous driving, where it enhances navigation and decision-making. However, most existing SLAM systems are heavily reliant on the assumption of a static world, including state-of-the-art (SOTA) systems such as ORB-SLAM2 [ORB-SLAM2], ORB-SLAM3 [ORB-SLAM3], LSD-SLAM [LSD-SLAM] and Direct Sparse Odometry [engel2017direct]. These systems typically perform with high precision in environments with predominantly static objects. Nevertheless, their precision tends to deteriorate when deployed in dynamic environments, which are common in real world scenarios. Some algorithms perform well in dynamic scenes with the utilization of robust kernels like Huber kernel [chebrolu2021adaptive], but it relies heavily on the accuracy of the depth measurement obtained from LiDAR sensor. Research efforts aimed at addressing the degradation of SLAM systems in dynamic environments can be categorized into three types. Type I focuses only on geometric methods, which involve classifying feature points based solely on geometric information. This information includes epipolar error, reprojection error, image intensity, depth, and optical flow, as demonstrated in works such as [zhu2022nice]. Type II, exemplified by works like DynaSLAM [DynaSLAM], leverages prior class information obtained from object detection models or semantic segmentation models to identify features within bounding boxes or masks of objects belonging to dynamic classes as outliers, which are then discarded during camera pose estimation. While these approaches partially mitigate the degradation of SLAM systems in dynamic environments, their effectiveness is limited. For instance, features extracted from stationary parked cars on the roadside may be erroneously removed, despite their static nature and potential value in estimating camera pose, leading to information loss. In Figure 1, a qualitative comparison is presented between our Multilayer Perceptron (MLP)-based discriminator for dynamic and static feature points and the method proposed in DynaSLAM, using a snapshot from the KITTI Odometry dataset [geiger2013visionKITTIDatasets]. The top image displays the results of our method, while the bottom image shows the results of the method in DynaSLAM. In the left image, the method in DynaSLAM categorizes feature points on the highway guardrail as dynamic, whereas our method correctly predicts them as static. In the right image, numerous static feature points associated with parked cars on the roadside are misclassified as dynamic by the method in DynaSLAM, whereas our method correctly identifies them as static. This evidence supports the effectiveness and reliability of our method. Type III represents the current mainstream approach, which combines methods from Type I and Type II, as demonstrated in works such as DynaSLAM II [DynaSLAMII], CFP-SLAM [CFP-SLAM], PointSLOT [PointSLOT], VDO-SLAM [VDO-SLAM], and D-SLAM [wen2023dynamic], by initially acquiring prior class information through deep learning models such as object detection models or semantic segmentation models. The YOLO series is the most common deep learning model in the SLAM domain, known for its high precision and speed, such as YOLOv5 [jocher2022ultralytics] and YOLOACT [bolya2019yolact]. Subsequently, geometric methods are used to eliminate dynamic feature points by leveraging prior model information. These methods consider factors such as epipolar error, re-projection error, image intensity, depth, and velocity. However, many methodologies in Type III rely on only partial geometric information. For instance, some methods solely focus on epipolar error [yu2018ds], re-projection error [PointSLOT], or a combination thereof [CFP-SLAM]. These approaches are commonly employed in the dynamic SLAM field. Nevertheless, these methods often require manual parameter tuning, lack generality, and are excessively laborious to fine-tune. Furthermore, the evaluation of methods distinguishing between dynamic and static feature points is typically conducted indirectly through SLAM field metrics, as there is a lack of publicly available datasets specifically designed for this classification task. This indirect evaluation method tends to be inaccurate and unreliable due to the randomness in the Random Sample Consensus (RANSAC) process in camera pose estimation within the SLAM system.
@@
To address the aforementioned gaps, we introduce a MLP-based real-time stereo SLAM system and establishes publicly available datasets for evaluating dynamic and static feature points method directly. The contributions outlined in this work are as follows:
@@
  • We introduce an open-source MLP-based real-time stereo system designed to enhance performance in dynamic environments. This system leverages comprehensive geometry information and eliminates the need for manual parameter configuration. To the best of our knowledge, this is the first time a machine learning-based method has been introduced to distinguish dynamic and static feature points in the SLAM field.
@@
  • We have constructed publicly available datasets comprising more than 50,000 feature points sourced from the KITTI Odometry datasets [geiger2013visionKITTIDatasets] for the purpose of training and testing dynamic and static feature points discriminators. The efficacy of it can be evaluated directly with this dataset.
  • We elaborate our methods ’s performance in all.
@@ II-A Dynamic SLAM
DynaSLAM has gained significant attention in Visual Simultaneous Localization and Mapping since its release. It identifies feature points within the semantic masks of potential moving objects as outliers and discards them. Although it improves performance in the KITTI Odometry dataset, it may misrepresent dynamic and static objects, as shown in Figure 1. Furthermore, DynaSLAM experiences high latency due to its dependence on the computationally intensive MaskRCNN model [he2017mask] for semantic segmentation. Subsequent research endeavors have aimed to address the degradation of SLAM systems in dynamic environments. For instance, work CFP-SLAM[PointSLOT] introduced a Moving Objects Recognition (MOR) approach to differentiate between dynamic and static feature points by utilizing the re-projection error. In comparison, work [CFP-SLAM] employs a DBSCAN[ester1996densityDBSCAN] clustering machine learning technique to filter out features within bounding boxes that belong to the background and then employ a geometric method that incorporates geometric information such as re-projection error and epipolar error to distinguish dynamic and static feature points. An inherent limitation in [PointSLOT] and [CFP-SLAM] is the requirement of the establishment and tuning of manual parameters, which restricts the application scenarios and diminishes the generality of the SLAM system. Furthermore, these approaches do not fully leverage geometric information. The VDO-SLAM[VDO-SLAM] and DynaSLAM II[DynaSLAMII] are are extensions of DynaSLAM, with DynaSLAM II representing the current state-of-the-art (SOTA) in camera motion estimation. Both works initially estimate the camera pose using feature points outside the potential dynamic class semantic mask, akin to the approach in DynaSLAM. Subsequently, they estimate the object pose for each frame and optimize the camera and object poses simultaneously through Bundle Adjustment (BA) factor graph. However, these methodologies encounter high latency issues due to data association challenges between dynamic objects and object pose estimation. Additionally, the precision of object pose estimation is compromised by the limited number of matched feature points and the quality of feature matching for each object.
@@ II-B Multilayer Perceptron
Multilayer Perceptron (MLP) [rosenblatt1961principles] is commonly utilized as a tool for classification tasks, and it has been effectively employed in artificial intelligence, particularly in the field of computer vision. Some renowned applications include the utilization of MLP to classify 3d point cloud models [qi2017pointnet] and for natural language processing [vaswani2017attention]. In the field of SLAM, studies [sucar2021imap, zhu2022nice] apply the MLP to decode grid features into occupancy values in the mapping part, study [mildenhall2021nerf] uses the MLP to implicitly represent feature points in space. To the best of our knowledge, the method introduced in our letter is the first instance of applying the MLP as a discriminator for dynamic and static feature points in the SLAM domain.
@@ III-A Notations in SLAM
World and camera coordinates are denoted by {WW} and {CC} respectively. The $i_{\text{th}}$ feature point is denoted by $x_{i}$, and its depth is represented by $d_{i}$. The coordinate of feature point $x_{i}$ in camera and world coordinate are describe by $^{t}X^{B}_{i}$, $B\in\{W,C\}$. Moreover, $^{t}P_{i}$ marks the location of feature point $x_{i}$ in the image plane, and the operation $\pi(.)$ represents projecting an 3D point cloud from camera coordinates into the image plane. i.e. $^{t}P_{i}=\pi(^{t}X^{C}_{i})$. Besides, the translation matrix $^{t}T_{CW}=\begin{bmatrix}{}^{t}R_{CW},\textbf{t}_{CW}\\ 0,1\end{bmatrix}\in SE(3)$ is used to describe a transformation from the world coordinate $\{W\}$ to the camera coordinate $\{C\}$. i.e. $^{t}X^{C}_{i}={{}^{t}R_{CW}}{{}^{t}X^{W}_{i}}+{t}_{CW}$ . The essential matrix between t1t-1 frame and $t$ frame is denoted by $^{t}E$.
@@ III-B Camera Pose Estimation
In the camera pose estimation process, we search the static feature point correspondences among consecutive frames to avoid the impact of dynamic objects. The camera pose is acquired through minimizing the re-projection error of static feature point correspondences, which is described by:
@@
$^{t}e_{i}={{}^{t}}P_{i}-\pi(^{t}T_{CW}{{{}^{t-1}}\hat{X}^{W}_{i}}),$ (2)
and the operation $\rho(\cdot)$ is a robust kernel function for reducing impact of feature points with big error, and $N$ denotes the number of static feature correspondences. The $^{t-1}\hat{X}^{W}_{i}$ indicates homogeneous coordinate of $^{t-1}X^{W}_{i}$}.
@@ III-C Notations in MLP
$e_{I}({{}^{t-1}x_{i}},^{t}x_{i})$ $=\left\lVert I({{}^{t-1}P_{i}})-I({{}^{t}P_{i}})\right\rVert^{2}_{2},$ (3)
$e_{D}(^{t}E,{{}^{t-1}x_{i}},^{t}x_{i})$ $=\frac{{{{}^{t}P_{i}}}^{T}{{}^{t}E}}{{\left\lVert{{{}^{t}P_{i}}}{{}^{t}E}\right\rVert_{2}}}\cdot{{}^{t-1}P_{i}},$ (4)
$e_{Re}(R,t,{{}^{t-1}x_{i}},^{t}x_{i})$ $=\left\lVert\pi(R\pi({{}^{t}P_{i}},d_{i})^{-1}+t)-{{}^{t-1}P_{i}}\right\rVert^{2}_{2}.$ (5)
We use operation $M(\cdot)$ to denote applying MLP model for classifying dynamic and static feature points. The prediction result denoted by $y_{i}\in R^{2}$. And the class of each feature point is denoted by $c_{i}$. For our problem, the class $c_{i}$ equals 0 or $1$. We used 0 to indicate static class and $1$ for dynamic class. Thus, the MLP model predicted process can be described by $c_{i}=argmax(y_{i})$, where $y_{i}=M(v_{i})$.
III-D System Structure
The diagram of our MLP-based stereo real-time SLAM system is shown in Figure 2. It is built on ORB-SLAM2 with an extra dynamic and static object discriminator module to filter dynamic feature points. Initially, the input stereo images undergo ORB feature extraction. Subsequently, object detection and tracking are performed on the left image. The process then moves to a coarse estimation stage where features outside the bounding box of potential dynamic objects, and the background static features within the bounding box that are filtered by the Depth Filter Module (DFM), are used for coarse camera pose estimation. The DFM will be explained in detail in Section IV-A. Then each feature point $x_{i}$ that is inside the bounding box and not filtered by the DFM generates an input feature $v_{i}$ using Eq. (3), (4), and (5). An example of the bounding box is shown in Figure 2 as the green box around the car. The MLP model assigns a label $c_{i}$ to each feature point to distinguish static and dynamic feature points. Finally, all static feature points are used for fine camera pose estimation. Feature point flows across the different modules in our method are shown in Figure 3 .
@@ IV Proposed Method
Figure 2: Diagram of our proposed MLP based real-time stereo SLAM system. It was built on ORB-SLAM2. There are four main parts in total. The blue section involves pre-processing, feature extraction, object detection, and object tracking. The pre-processing and feature extraction parts are default setting of ORB-SLAM2.The green section is for coarse estimation, generating a rough pose estimation. The pink section involves an MLP model for classifying dynamic and static feature points within potential dynamic objects. The yellow section is for fine estimation, refining the camera pose generated in the coarse estimation stage by minimizing the re-projection error of all static feature points. These points include those outside the bounding box of potential dynamic objects, filtered by the Depth Filter Module, and classified as static by the MLP model.
Figure 3: Feature points flow among different module of our SLAM system. The yellow circle represented dynamic feature points set, and the blue one standed for static feature points set. Moreover, all square represented an module of our SLAM system, the one outline with dashed line indicated that it has no impact to feature point sets.
In this section, details of each module and our collected dynamic and static feature points classification dataset will be elaborated.
IV-A Depth Filter Module (DFM)
We opt for object detection and tracking models over semantic segmentation models in our SLAM system to reduce latency, as the latter typically requires more computational resources. However, object detection models produce less accurate bounding boxes compared to mask from semantic segmentation models. To address this issue, we introduce the DFM to remove background features within the bounding box, classifying them as static. The process of the DFM is as follows: Initially, we calculate the average depth $\mu_{d}$ of all feature points within the bounding box, provided there are a sufficient number of feature points. Subsequently, we calculate the standard deviation $\delta_{d}$ of these points. Any feature points falling outside the range $\mu_{d}-\eta\delta_{d},\mu_{d}+\eta\delta_{d}]$ are eliminated from the potential dynamic feature point set and treated as static in coarse estimation stage. Here, $\eta$ represents a constant value that describes the degree of deviation from the average depth value.